OpenStudy (lukecrayonz):

Solve the integral sqrt(x)-2 from [1,9]

OpenStudy (lukecrayonz):

\[\int\limits_{1}^{9}\sqrt(x)-2*dx\]

OpenStudy (abb0t):

\(\sf \color{red}{\sqrt{x} = x^{\frac{1}{2}}}\)

OpenStudy (anonymous):

\[\int_1^92dx=8\times 2\] is a start

OpenStudy (abb0t):

Then use the fundamental theorem of calculus pt. II to solve. Which is F(b)-F(a)

OpenStudy (shamil98):

what, lord abb0t said.

OpenStudy (lukecrayonz):

 My face when you said that

OpenStudy (lukecrayonz):

Kk let me quickly do the work.

OpenStudy (abb0t):

chop-chop!

OpenStudy (lukecrayonz):

(4/3)?

OpenStudy (shamil98):

\[\large \int\limits_{1}^{9} x^{\frac{ 1 }{ 2} } - 2~dx\] \[\Huge \int\limits_{1}^{9} \frac{ x^\frac{ 3 }{ 2 } }{ \frac{ 3 }{ 2 } } - 2x\]

OpenStudy (shamil98):

\[\Huge \int\limits_{1}^{9} \frac{ 9^\frac{ 3 }{ 2 } }{ \frac{ 3 }{ 2 } } - 2(9) - (\frac{ 1^{\frac{ 3 }{ 2 }} }{ \frac{ 3 }{ 2 } } - 2(1))\]